OpenStudy (anonymous):

I need help with absolute value 2[z+1]-3=z-2 Check for extraneous solutions.

OpenStudy (anonymous):

absolute value is just the numbers distance from zero..... |-17| this number is 17 from zero so its absolute value is 17 |4| = 4 |-8| = 8 |34| = 34

OpenStudy (anonymous):

You have to solve it like youre solving an equation though..

OpenStudy (anonymous):

z-1= 2*1=2 -3=-2 =-4

OpenStudy (anonymous):

That doesn't make sense...

OpenStudy (anonymous):

yes it does hold up its -1

OpenStudy (anonymous):

you solved it like an equation and got -1?

OpenStudy (anonymous):

I got 0

OpenStudy (anonymous):

what